# Instructions for ACL 2023 Proceedings

### v1

@@ Abstract
Web-scale training on paired text-image data is becoming increasingly central to multimodal learning, but is challenged by the highly noisy nature of datasets in the wild. Standard data filtering approaches succeed in removing mismatched text-image pairs, but permit semantically related but highly abstract or subjective text. These approaches lack the fine-grained ability to isolate the most concrete samples that provide the strongest signal for learning in a noisy dataset. In this work, we propose a new metric, Image Caption Concreteness (ICC), that evaluates caption text without an image reference to measure its concreteness and relevancy for use in multimodal learning. Our approach leverages strong foundation models for measuring visual-semantic information loss in multimodal representations. We demonstrate that this strongly correlates with human evaluation of concreteness in both single-word and sentence-level texts. Moreover, we show that curation using ICC complements existing approaches: It succeeds in selecting the highest quality samples from multimodal web-scale datasets to allow for efficient training in resource-constrained settings.
@@ 1 Introduction
$\uparrow$
A sandwich sits on a small blue plate (0.96)
Curly-haired man with a mustache in a vintage photo (1.0)
A cat standing on a counter looking at a coffee cup (1.0)
$\downarrow$
It does not look like something I would eat (0.02)
Talk about a bad hair day, his is frightful (0.01)
I cant see this image it is too dark (0.02)
@@
Existing datasets are commonly filtered using VLMs such as CLIP Radford et al. (2021) to identify image-text semantic misalignments Sharma et al. (2018); Schuhmann et al. (2022), i.e. captions irrelevant to their images; using rule-based proxies such as measuring the complexity of captions via semantic parsing Radenovic et al. (2023); or removing images that contain text that overlaps with the caption Maini et al. (2023). However, these approaches fail to identify captions that are highly abstract and may contain subjective, non-visual information, despite being semantically aligned with the image and having a sufficiently complex grammar. Figure 1 shows examples of such image-caption pairs. A caption such as “It does not look like something I would want to eat” is semantically related to the image, yielding high CLIP similarity, but contains subjective details which provide a confounding signal when training VLMs (See also Figure 2). A model trained to generate such captions from images may learn to hallucinate details, e.g., liking a certain type of food in our example, which are not visually grounded and are highly subjective. Similarly, such image-caption pairs provide a weaker signal for representation learning than images with visually-concrete captions (e.g. “A sandwich sits on a small blue plate”), which may impede the learning process – particularly in a resource-restricted setting where data or compute is limited.
@@
Thus, we suggest filtering image captions by their visual concreteness, i.e., the degree to which a text describing a specific visual scene can be vividly imagined (unlike abstract text that may correspond to many possible visual representations or include non-visual contextual information). We show that this new dimension of textual quality enables selecting image-caption pairs that provide a strong supervision signal for vision-and-language tasks, particularly in resource-constrained settings where training directly on noisy web-scale multimodal data fails to converge to a satisfactory solution in a limited number of iterations.
@@
We propose the Image Caption Concreteness (ICC) metric for quantifying the visual concreteness of image captions calculated from text alone, i.e., without an image reference. We measure concreteness using autoencoding pipelines with visual-semantic information bottlenecks. Specifically, we use a visual-bottleneck autoencoder that leverages text-to-image generative models’ competence and a semantic-bottleneck autoencoder that identifies how well a large language model (LLM) recovers the input caption from its semantic CLIP embedding. As these models require costly inference through large generative models, they cannot feasibly run on a large scale; therefore, our ICC metric is distilled from these pipelines, enabling fast, computationally-efficient inference.
@@
Stated explicitly, our contributions are as follows: (1) We propose the ICC metric distilled from foundation VLM models with a novel combination of autoencoding pipelines; (2) we show that ICC is highly correlated to human concreteness judgements of caption texts; (3) we demonstrate that ICC succeeds in selecting a core of samples from web-scale image-caption datasets for vision-and-language tasks, with superior downstream performance to existing multimodal filtering methods; (4) we will release our data, code, and models, anticipating the use of ICC for further tasks that require curation of high-quality image-caption pairs from noisy web-scale multimodal data.
Refer to caption (a) poaching still remains the biggest threat to tigers Refer to caption (b) Wheat bread is always the healthy choice for lunchtime
Refer to caption (c) want a ring like this! Refer to caption (d) Someone did not observe the stop sign and now it is knocked over
@@ 2 Method
As an example, consider the text “Wheat bread is always the healthy choice for lunchtime” in Figure 1(b). The notion of wheat bread being a healthy choice is inherently non-visual, and is unlikely to be directly encoded in a visual representation, e.g., in an image. Therefore, this information is likely to be lost in an autoencoding process that includes an image as the bottleneck, when the encoded image is decoded back to the textual modality.
@@
We proceed to describe our proposed visual-bottleneck autoencoder and semantic-bottleneck autoencoder components, and their consolidated distillation into the final ICC metric. See Figure 3 for an overview of our full pipeline.
@@
Visual-Bottleneck Autoencoder (VBA). Since a caption represents an image, we construct the VBA by using an image as an intermediate representation via which textual information passes. In particular, we concatenate a text-to-image model (Stable Diffusion 2, Ramesh et al., 2022) and a captioning model (BLIP-2, Li et al., 2023) as shown in Figure 3 (bottom left). This autoencoding pipeline measures text concreteness by encoding and decoding a caption, followed by measuring semantic fidelity in reconstruction using BERTScore (F1) Zhang et al. (2019).
@@
While the VBA pipeline is a simple and intuitive way of enforcing a visual bottleneck, it may sometimes produce sub-optimal reconstructions even for highly visual texts due to its inherently lossy nature. For example, the caption “a small black french bulldog” in Figure 3 may be reconstructed by the VBA from the generated image to “a dog with a white chest", which is relatively semantically far from the original caption and thus results in a relatively low reconstruction score of 0.6 for a concrete caption. This stems from the dense information content of generated images, which may contain details (such as the dog’s white chest) which were not mentioned explicitly in the original caption, and from the tendency of the captioning decoder to focus on different details than those used to generate the image. To alleviate this issue, we suggest next a complementary method which, while still maintaining a visual prior, also enforces a strong prior on the caption semantic information.
@@
Semantic-bottleneck Autoencoder (SBA). Motivated by findings that CLIP embeddings encode visual information in text and particularly concreteness Alper et al. (2023), we construct an autoencoding pipeline with CLIP text embeddings as a semantic information bottleneck, as shown in Figure 3 (top left). We extract visual information from the CLIP text embedding space by utilizing a frozen LLM (Llama-2-7b, Touvron et al., 2023), by training a linear layer that converts CLIP text encoder’s output to inputs for the LLM. This enables more efficient training compared to fine-tuning the LLM. The training objective aims at reconstructing the input captions via a token-wise cross-entropy objective.
@@
After training SBA over image–caption pairs, we use it for measuring text concreteness by encoding and decoding the text followed by measuring reconstruction fidelity. To measure preservation of fine-grained textual details, we quantify this fidelity via per-character edit distance Levenshtein et al. (1966), standardized by caption length, as detailed in Appendix A.1.
@@
This pipeline generally succeeds in reconstructing highly concrete text (such as “A small black french bulldog” shown in the top left part of Figure 3). However, the strong textual prior of the SBA may also leak information about abstract and subjective captions as well (e.g. the abstract caption “A nice location” yields a relatively high reconstruction score of 0.4), limiting its correlation with visual concreteness. Overall, the SBA and VBA provide complementary scores, where each correlates more strongly to visual concreteness in different cases. Therefore, they perform most strongly when combined together, as we explicitly verify in our ablations in Section 4. We also show qualitative examples in figures 9 and 10 in the appendix.
Optimal Score Combination. To compute the optimal combination of the two scores, we label 244 captions, sampled uniformly over VBA and SBA scores, with concreteness scores in the range 003333. We use logistic regression to find the parameters $a,b,c$ of $\sigma(a\cdot VBA{}+b\cdot SBA{}+c)$, where $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function, such that the output will approach 1 for concrete captions and 0 for abstract ones. We label concrete captions as captions with concreteness above the median score in the labeled dataset and abstract captions as captions with a score below this median. We visualize the annotated samples and the regression line $a\cdot VBA{}+b\cdot SBA{}+c=0$ in Figure 4. As seen in the figure, both scores contribute to the optimal predicted concreteness score, validating the importance of using both SBA and VBA components together in our full pipeline.
@@
ICC Distillation. Using the aforementioned pipelines to quantify the concreteness at scale is not feasible, as this requires running large models (e.g., diffusion models, LLMs) with billions of parameters for many forward passes per instance (up to dozens of forward passes for the diffusion models inference and for the LLM and captioning model decoding). This requires more than 1,000 GPU hours for a dataset of 1M samples. Therefore, we assemble SBA and VBA reconstruction scores over a relatively small collection of image-caption pairs and distill their aggregated values into our final ICC score. This enables efficient inference that can easily run on a large scale, with over a hundred times faster inference time and much less compute required. Specifically, we train a small text encoder model Liumm et al. (2019) over the optimal combination of the two score.
@@
Implementation Details of ICC Construction. For the construction of our ICC score, we use a subset of CC3M Sharma et al. (2018) composed of 595K image-caption pairs, introduced by Liu et al. (2023) and designed to have wider concept coverage. We take a subset of 476K samples for training the linear layer of the SBA, and train for 2 epochs with a batch size of 128 and learning rate of 2e-3 with cosine scheduling function. The remaining 118K samples are used for generating reconstruction scores through the VBA and the trained SBA. For each input caption, we generate five reconstructed captions using beam search (five beams) with the VBA’s captioner and the SBA’s LLM and then choose the reconstructed caption with the highest similarity to the source caption. By generating the reconstructions and measuring the reconstruction fidelities, we obtain a dataset of 118K captions and corresponding reconstruction scores. We standardize by caption length to disentangle the dependency of the reconstruction scores to the caption length (i.e., forcing the same distribution of scores for all caption lengths), as described in the appendix, and train a small language model (DistillRoberta-Base) over the optimal combination of the two scores as previously described, over these 118K samples with a Mean Squared Error objective. This final distilled model is used for generating the ICC scores.
@@ 3.1 VLM Dataset Curation
Experimental Settings We investigate the effect of ICC and other filtering methods for curating a core of high-quality image-caption pairs from large multimodal datasets, comparing their effects on downstream task performance – both discriminative (representation learning) and generative (image captioning). We follow similar settings as described in the Datacomp Gadre et al. (2023) benchmark’s filtering track [^1] As opposed to the BYOD track which allows for modifying the samples, for instance by using synthetic captions., with the following modifications to model the resource-limited setting: given a training dataset comprised of $\mathcal{M}$ samples, the downstream model is constrained to train for exactly $\mathcal{N}\ll\mathcal{M}$ iterations over the filtered subset of the dataset. This contrasts with the original Datacomp setting where $\mathcal{N}=\mathcal{M}$, which requires significant compute for a web-scale dataset. Our formulation tests the ability of filtering methods to curate high-quality core subsets of such datasets. Our initial subset of LAION-400M is composed of $\mathcal{M}=8$ million samples and we fix $\mathcal{N}=2$ million training iterations. To verify the robustness of our method, we measure downstream performance over visually grounded benchmarks across three different sizes of filtering.
@@
We compare to three existing filtering methods – CLIPScore  Hessel et al. (2021), Complexity and Action (CA) Radenovic et al. (2023) using our re-implementation as there is no publicly-available code, and T-MARS Maini et al. (2023). CA is a rule-based filtering method which aims to retain only sufficiently complex captions that also contain an action, based on semantic parsing. T-MARS filters multimodal datasets by removing samples whenever an image includes text that overlaps significantly with the caption. As opposed to these method, we focus on filtering according to the level of concreteness of image captions.
CA
Bathroom and bathroom remodeling ideas…
How to make a flower bouquet
Buster Posey hits a two-run home run…
TMS
Wholesale Best Price For White Plastic Bathroom…
Coffee Tabletop Decorating Ideas
Cleveland Indians vs. Boston Red Sox
CLIP
Aquariums-Aquarium of the Year - AQUARIUM OF THE YEAR
Dandelion Flower Gift Ideas for Kids…
Bryce Harper of the Toronto Blue Jays signs autographs for fans prior to the game
ICC
White Plastic Toilet Seat
White and Yellow Gerbera Flowers in a Glass Vase
black and white photo of a baseball player
@@
Figure 5: Downsream captioning qualitative examples. Captions generated by the model trained over ICC’s filtered dataset generates more concrete and accurate captions compared to models trained over dataset filtered with CA, T-MARS (TMS) and CLIP. Images are from MS-COCO dataset.
Captioning Models.
@@
Method Dataset Size B@4 M R C S BSc B@4 M R C S BSc
Random 100k 0.9 4.7 11.2 5 2.3 0.64 1.0 5.1 11.9 5.6 1.6 0.68
CLIP 100k 1.1 5.5 11.9 2.5 2.2 0.75 1.4 5.7 12 3 1.5 0.71
CA 100k 0.9 3.7 7.3 3.2 1.6 0.27 1.6 4.4 9.4 4.1 1.2 0.33
T-MARS 100k 1.2 4.6 10.6 5.6 2.3 0.53 1.3 4.9 11.6 6.3 1.7 0.61
ICC 100k 10.1 15.4 35.4 35.8 10.3 0.9 12.1 15.8 35.9 33.3 6.4 0.9
Random 200k 0.9 4.2 9.8 5 2.2 0.51 1 4.8 11.2 5.9 1.7 0.6
CLIP 200k 1.3 5.7 12.4 3.4 2.6 0.72 1.6 6 12.6 3.5 1.8 0.67
CA 200k 0.5 2.8 5.7 3.7 1.3 0.18 1.2 3.4 7.2 4.1 1.1 0.24
T-MARS 200k 1.1 4.6 10.7 6.5 2.4 0.5 1.7 5.4 12.3 7.8 1.9 0.6
ICC 200k 10.0 15.2 34.6 35.5 10.4 0.9 13.1 15.8 35.2 34.3 6.7 0.9
Random 500k 0.6 3.4 8.0 4.5 1.9 0.42 0.9 4.2 10.1 5.5 1.5 0.55
CLIP 500k 5.2 9.4 22 15.1 5.3 0.8 5.2 8.9 21.3 12.9 3 0.8
CA 500k 0.7 3.1 6 3.6 1.4 0.19 2.1 4.5 9.4 5.3 1.5 0.29
T-MARS 500k 0.8 3.7 8.9 5.7 2 0.42 1.2 4.7 10.8 6.5 1.7 0.65
ICC 500k 8.3 13.9 31.4 30.9 9.7 0.89 10 14.2 31.3 28.2 6 0.89
@@
Table 1: Captioning results for different filtered dataset sizes. We perform evaluation of captioning models over MS-COCO and NoCaps datasets trained over different filtering schemes of the LAION-400M dataset, with varying dataset sizes. We compare the performance of ICC to three filtering baselines; CLIP indicates filtering by top CLIPScore, CA indicates Complexity and Action filtering and Random refers to random samples from LAION-400M. B@4, M, R, C, S and BSc denote BLEU-4, METEOR, Rouge-L, CIDEr, SPICE, and BERTScore metrics respectively. Best results are in bold.
@@
In Table 1 we show quantitative results of applying ICC filtering on top of standard CLIP filtering over the subset of LAION-400M for training a captioning model. The captioning model used is an encoder-decoder architecture with a pretrained Swin Liu et al. (2021) vision encoder and GPT-2 Radford et al. (2019) text decoder. We use a batch size of 100, and learning rate of 2e-5 with a cosine scheduler. We test our approach over two standard captioning benchmarks datasets – MS-COCO Lin et al. (2014) and NoCaps  Agrawal et al. (2019), across multiple captioning metrics. As illustrated in the table, filtering with ICC outperforms by a large margin the alternative filtering methods for captioning given a fixed number of desired samples and training iterations. Note that unlike other methods, ICC is directly aligned with the captioning objective, as a captioning model should generate visually-grounded concrete text. This can explain the large gap in performance between ICC and other filtering baselines. We show qualitative comparison between captioning models trained with different filtering methods in Figure 5, exemplifying how filtering with ICC promotes more concrete and accurate captioning.
Image-Text Representation Learning.

### v4

# ICC: Quantifying Image Caption Concreteness for Multimodal Dataset Curation

@@ Abstract
Web-scale training on paired text-image data is becoming increasingly central to multimodal learning, but is challenged by the highly noisy nature of datasets in the wild. Standard data filtering approaches succeed in removing mismatched text-image pairs, but permit semantically related but highly abstract or subjective text. These approaches lack the fine-grained ability to isolate the most concrete samples that provide the strongest signal for learning in a noisy dataset. In this work, we propose a new metric, Image Caption Concreteness (ICC), that evaluates caption text without an image reference to measure its concreteness and relevancy for use in multimodal learning. Our unsupervised approach leverages strong foundation models for measuring visual-semantic information loss in multimodal representations. We demonstrate that this strongly correlates with human evaluation of concreteness in both single-word and caption-level texts. Moreover, we show that curation using ICC complements existing approaches: It succeeds in selecting the highest quality samples from multimodal web-scale datasets to allow for efficient training in resource-constrained settings.
@@ 1 Introduction
Refer to caption Refer to caption Refer to caption
$\uparrow$ A sandwich sits on a small blue plate (0.96) Curly-haired man with a mustache in a vintage photo (1.0) A cat standing on a counter looking at a coffee cup (1.0)
$\downarrow$ It does not look like something I would eat (0.02) Talk about a bad hair day, his is frightful (0.01) I cant see this image it is too dark (0.02)
@@
Existing datasets are commonly filtered using VLMs such as CLIP Radford et al. (2021) to identify image-text semantic misalignments Sharma et al. (2018); Schuhmann et al. (2022), i.e. captions irrelevant to their images; using rule-based proxies such as measuring the complexity of captions via semantic parsing Radenovic et al. (2023); or removing images that contain text that overlaps with the caption Maini et al. (2023). However, these approaches fail to identify captions that are highly abstract and may contain subjective, non-visual information, despite being semantically aligned with the image and having a sufficiently complex grammar. Figure 1 shows examples of such image-caption pairs. A caption such as “It does not look like something I would want to eat” is semantically related to the image, yielding high CLIP similarity, but contains subjective details which provide a confounding signal when training VLMs (See also Figure 2). A model trained to generate such captions from images may learn to hallucinate details, e.g., liking a certain type of food in our example, which are not visually grounded and are highly subjective. Similarly, such image-caption pairs provide a weaker signal for representation learning than images with visually concrete captions (e.g. “A sandwich sits on a small blue plate”), which may impede the learning process – particularly in a resource-restricted setting where data or compute is limited.
@@
Thus, we suggest filtering image captions by their visual concreteness, referring to the extent to which a text describes visual aspects of a scene in a manner that can be vividly imagined Schwanenflugel (2013); Hessel et al. (2018) [^1] Some works have treated this as roughly synonymous with imageability (visual association), while others use concreteness to refer more generally to association with sensory experiences of all types Richardson (1975); Khanna and Cortese (2021). Our work focuses on the visual modality.. This contrasts with abstract text, which may correspond to many possible visual interpretations or include subjective information. We show that this new dimension of textual quality enables selecting image-caption pairs that provide a strong supervision signal for vision-and-language tasks, particularly in resource-constrained settings where training directly on noisy web-scale multimodal data fails to converge to a satisfactory solution in a limited number of iterations.
@@
We propose the Image Caption Concreteness (ICC) metric for quantifying the visual concreteness of image captions calculated from text alone, i.e., without an image reference. We measure concreteness using unsupervised autoencoding pipelines with visual-semantic information bottlenecks. Specifically, we use a visual-bottleneck autoencoder that leverages text-to-image generative models’ competence and a semantic-bottleneck autoencoder that identifies how well a large language model (LLM) recovers the input caption from its semantic CLIP embedding. As these models require costly inference through large generative models, they cannot feasibly run on a large scale; therefore, our ICC metric is distilled from these pipelines, enabling fast, computationally-efficient inference.
@@
Stated explicitly, our contributions are as follows: (1) We propose the ICC metric distilled from foundation VLM models with a novel combination of unsupervised autoencoding pipelines; (2) we show that ICC highly correlates to human concreteness judgements of caption texts; (3) we demonstrate that ICC succeeds in selecting a core of samples from web-scale image-caption datasets for vision-and-language tasks, with superior downstream performance to existing filtering methods.
(a) poaching still remains the biggest threat to tigers
(b) Wheat bread is always the healthy choice for lunchtime
(c) want a ring like this!
(d) Someone did not observe the stop sign and now it is knocked over
@@ 2 Method
As an example, consider the text “Wheat bread is always the healthy choice for lunchtime” in Figure 2(b). The notion of wheat bread being a healthy choice is inherently non-visual and is unlikely to be directly depicted in an image. Therefore, this information is likely to be lost in an autoencoding process that includes an image as the bottleneck, when the encoded image is decoded back to the textual modality.
@@
We proceed to describe our proposed visual-bottleneck autoencoder and semantic-bottleneck autoencoder components, and their distillation into the final ICC metric. See Figure 3 for an overview of our full pipeline.
@@
Visual-Bottleneck Autoencoder (VBA). Since a caption represents an image, we construct the VBA by using an image as an intermediate representation via which textual information passes. In particular, we concatenate a frozen text-to-image model (Stable Diffusion 2, Ramesh et al., 2022) and a frozen captioning model (BLIP-2, Li et al., 2023) as shown in Figure 3 (bottom left). This autoencoding pipeline measures text concreteness by encoding and decoding a caption, followed by measuring semantic fidelity in reconstruction using BERTScore (F1) Zhang et al. (2019). We note that this pipeline contains no trained parameters as it concatenates pretrained, frozen models.
@@
While the VBA pipeline is a simple and intuitive way of enforcing a visual bottleneck, it may sometimes produce sub-optimal reconstructions even for highly visual texts due to its inherently lossy nature. For example, the caption “a small black french bulldog” in Figure 3 may be reconstructed by the VBA from the generated image to “a dog with a white chest", which is relatively semantically far from the original caption and thus results in a relatively low reconstruction score of 0.6 for a concrete caption. This stems from the dense information content of generated images, which may contain details (such as the dog’s white chest) which were not mentioned explicitly in the original caption, and from the tendency of the captioning decoder to focus on different details than those used to generate the image. To alleviate this issue, we proceed to propose a complementary method using a stronger prior on caption semantics.
@@
Semantic-bottleneck Autoencoder (SBA). Motivated by findings that CLIP embeddings encode visual information in text and particularly concreteness Alper et al. (2023), we construct an autoencoding pipeline with CLIP text embeddings as a semantic information bottleneck, as shown in Figure 3 (top left). We extract visual information from the CLIP text embedding space by utilizing a frozen LLM (Llama-2-7b, Touvron et al., 2023), by training a linear layer that converts CLIP text encoder’s output to inputs for the LLM. The training objective aims at reconstructing the input captions via a token-wise cross-entropy objective. By keeping the encoder backbone (CLIP) frozen, this introduces an information bottleneck preventing faithful reconstruction of abstract texts.
@@
After training the SBA over image–caption pairs, we use it for measuring text concreteness by encoding and decoding the text followed by measuring reconstruction fidelity. To measure preservation of fine-grained textual details, we quantify this fidelity via per-character edit distance Levenshtein et al. (1966), standardized by caption length, as detailed in Appendix A.1.
@@
This pipeline generally succeeds in reconstructing highly concrete text (such as “A small black french bulldog” shown in the top left part of Figure 3). However, the strong textual prior of the SBA may also leak information about abstract and subjective captions as well (e.g. the abstract caption “A nice location” yields a relatively high reconstruction score of 0.4), limiting its correlation with visual concreteness. Overall, the SBA and VBA provide complementary scores, where each correlates more strongly to visual concreteness in different cases. Therefore, they perform most strongly when combined together, as we explicitly verify in our ablations in Section 4. We also show qualitative examples in figures 8 and 9 in the appendix.
@@
ICC Distillation. Using the aforementioned pipelines to quantify the concreteness at scale is not feasible, as this requires running large models (e.g., diffusion models, LLMs) with billions of parameters for many forward passes per instance (up to dozens of forward passes for the diffusion models inference and for the LLM and captioning model decoding). This requires more than 1,000 GPU hours for a dataset of 1M samples. Therefore, we assemble SBA and VBA reconstruction scores over a relatively small collection of image-caption pairs and distill their aggregated values into our final ICC score. This enables efficient inference that can easily run on a large scale, with over a hundred times faster inference time and much less compute required. Specifically, we train a small text encoder model Liumm et al. (2019) to predict a logit-linear combination of the SBA and VBA scores, computed as described in the appendix.
@@
Implementation Details of ICC Construction. For the construction of our ICC score, we use a subset of CC3M Sharma et al. (2018) composed of 595K image-caption pairs, introduced by Liu et al. (2023) and designed to have wider concept coverage. We take a subset of 476K samples for training the linear layer of the SBA, and train for 2 epochs with a batch size of 128 and learning rate of 2e-3 with cosine scheduling function. The remaining 118K samples are used for generating reconstruction scores through the VBA and the trained SBA. For each input caption, we generate five reconstructed captions using beam search (five beams) with the VBA’s captioner and the SBA’s LLM and then choose the reconstructed caption with the highest similarity to the source caption. By generating the reconstructions and measuring the reconstruction fidelities, we obtain a dataset of 118K captions and corresponding reconstruction scores. We standardize by caption length to disentangle the dependency of the reconstruction scores to the caption length (i.e., forcing the same distribution of scores for all caption lengths), as described in the appendix. We train a small language model (DistillRoberta-Base) to predict the combined scores on these samples with a Mean Squared Error objective. This final distilled model is used for generating the ICC scores.
@@ 3.1 VLM Dataset Curation
Experimental Settings. We investigate the effect of ICC and other filtering methods for curating a core of high-quality image-caption pairs from large multimodal datasets, comparing their effects on downstream task performance – both discriminative (representation learning) and generative (image captioning). We follow similar settings as described in the Datacomp Gadre et al. (2023) benchmark’s filtering track [^2] As opposed to the BYOD track which allows for modifying the samples, for instance by using synthetic captions., with the following modifications to model the resource-limited setting: given a training dataset comprised of $\mathcal{M}$ samples, the downstream model is constrained to train for exactly $\mathcal{N}\ll\mathcal{M}$ iterations over the filtered subset of the dataset. This contrasts with the original Datacomp setting where $\mathcal{N}=\mathcal{M}$, which requires significant compute for a web-scale dataset. Our formulation tests the ability of filtering methods to curate high-quality core subsets of such datasets. Our initial subset of LAION-400M is composed of $\mathcal{M}=8M$ samples and we fix $\mathcal{N}=2M$ training iterations. To verify the robustness of our method, we measure downstream performance over visually grounded benchmarks across three different sizes of filtering.
@@
We compare to four existing filtering methods – CLIPScore  Hessel et al. (2021), Complexity and Action (CA) Radenovic et al. (2023) [^3] Using our re-implementation, as there is no publicly available code., T-MARS Maini et al. (2023), and PACScore Sarto et al. (2023). CA is a rule-based filtering method which aims to retain only sufficiently complex captions that also contain an action, based on semantic parsing. T-MARS filters multimodal datasets by removing samples whenever an image includes text that overlaps significantly with the caption. PACScore trains a CLIP-based model with positive-augmented contrastive learning approach, showing improved correlations with human intuition in scoring image-caption pairs. As opposed to these methods, we focus on filtering according to the concreteness of image captions.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption CA Tiger cubs playing in the rain at the Zoo of the Ozarks in Washington, D.C. on Saturday, Oct. 18 Coffee at the bar. I love this place! It’s a great way to get away from the hustle and bustle Cleveland Indians vs. Boston Red Sox Cambodia, the largest of all the African savannahs, is one of the most arid regions in the world. Rugby World Cup 2019: The men’s singles final takes place at the Ritz-Carlton in London, England, on Saturday, TMS Catching a lion in the wild is one of the most beautiful things you can do in the wild. Coffee at the bar. Photo credit: Flickr userfairy.com.au. (via Flickr) Buster Posey hits a two-run home run… Aerial view of a wild boar in a field in Namibia, South Africa, Africa. (Photo courtesy of the Namibian Wildlife) Astonishingly, there was no shortage of competition between the two-school teams at the London 2012 Olympic Games. CLIP Polar bear cubs pose for a photo with a polar bear in the background. Credit: NASA/JPL-Caltech/UCLA Pizza Hut Creamy Pizza Sandwich with Bacon, Cheese, and Tomato Sauce Bryce Harper of the Toronto Blue Jays signs autographs for fans prior to the game Aerial view of the world’s largest crocodile in the Serengeti National Park. New York Mets Fanatics Authentic 8"" x 10"" Skateboard Deck ICC Panda eating bamboo A picture of a pizza box full of pizzas black and white photo of a baseball player Zebra at the zoo A view of the tennis court from the front.
@@
Figure 4: Qualitative examples of captions generated by captioning models trained on datasets filtered with different filtering methods, over images from MS-COCO test split. CA denotes Complexity and Action filtering, and T-MARS is marked by TMS. As seen above, models trained on ICC-filtered data generate much more concrete and visually-grounded captions.
@@
Method # Samples B@4 M R C S BSc P B@4 M R C S BSc P
Random 100k 0.9 4.7 11.2 5 2.3 0.64 0.18 1 5.1 11.9 5.6 1.6 0.68 0.11
CLIP 100k 1.1 5.5 11.9 2.5 2.2 0.75 0.12 1.4 5.7 12 3 1.5 0.71 0.08
CA 100k 0.9 3.7 7.3 3.2 1.6 0.27 0.20 1.6 4.4 9.4 4.1 1.2 0.33 0.11
T-MARS 100k 1.2 4.6 10.6 5.6 2.3 0.53 0.20 1.3 4.9 11.6 6.3 1.7 0.61 0.11
PACScore 100k 1.9 6.8 15 5.5 3 0.82 0.16 2.9 7.5 16.1 7.7 2.4 0.79 0.1
ICC 100k 10.1 15.4 35.4 35.8 10.3 0.9 0.39 12.1 15.8 35.9 33.3 6.4 0.9 0.2
Random 200k 0.9 4.2 9.8 5 2.2 0.51 0.19 1 4.8 11.2 5.9 1.7 0.6 0.11
CLIP 200k 1.3 5.7 12.4 3.4 2.6 0.72 0.21 1.6 6 12.6 3.5 1.8 0.67 0.09
CA 200k 0.5 2.8 5.7 3.7 1.3 0.18 0.20 1.2 3.4 7.2 4.1 1.1 0.24 0.12
T-MARS 200k 1.1 4.6 10.7 6.5 2.4 0.5 0.21 1.7 5.4 12.3 7.8 1.9 0.6 0.12
PACScore 200k 2.9 7 15.5 7 3.7 0.73 0.19 3.8 7.4 15.8 8.8 2.6 0.67 0.12
ICC 200k 10 15.2 34.6 35.5 10.4 0.9 0.39 13.1 15.8 35.2 34.3 6.7 0.9 0.21
Random 500k 0.6 3.4 8 4.5 1.9 0.42 0.2 0.9 4.2 10.1 5.5 1.5 0.55 0.12
CLIP 500k 5.2 9.4 22 15.1 5.3 0.8 0.24 5.2 8.9 21.3 12.9 3 0.8 0.13
CA 500k 0.7 3.1 6 3.6 1.4 0.19 0.2 2.1 4.5 9.4 5.3 1.5 0.29 0.13
T-MARS 500k 0.8 3.7 8.9 5.7 2 0.42 0.21 1.2 4.7 10.8 6.5 1.7 0.65 0.12
PACScore 500k 2.6 6.5 15 8.6 3.7 0.65 0.21 3 6.9 15.4 10.4 2.6 0.65 0.13
ICC 500k 8.3 13.9 31.4 30.9 9.7 0.89 0.37 10 14.2 31.3 28.2 6 0.89 0.2
@@
Table 1: Captioning results for different filtered dataset sizes. We perform evaluation of captioning models over MS-COCO and NoCaps datasets trained over different filtering schemes of the LAION-400M dataset, with varying dataset sizes. We compare the performance of ICC to five filtering baselines. Among these, Random refers to random samples from LAION-400M, CLIP indicates filtering by top CLIPScore, and CA indicates Complexity and Action filtering. B@4, M, R, C, S, BSc and P denote BLEU-4, METEOR, Rouge-L, CIDEr, SPICE, BERTScore and Polos metrics respectively. # Samples denotes the amount of samples retained after filtering. Best results are in bold.
@@
Captioning Models. In Table 1 we show quantitative results of applying ICC filtering on top of standard CLIPScore filtering over the subset of LAION-400M for training a captioning model. The captioning model used is an encoder-decoder architecture with a pretrained Swin Liu et al. (2021) vision encoder and GPT-2 Radford et al. (2019) text decoder. We use a batch size of 100, and learning rate of 2e-5 with a cosine scheduler. We test our approach over two standard captioning benchmarks datasets – MS-COCO Lin et al. (2014) and NoCaps  Agrawal et al. (2019), across multiple captioning metrics Papineni et al. (2002); Banerjee and Lavie (2005); Lin (2004); Vedantam et al. (2015); Anderson et al. (2016); Zhang et al. (2019); Wada et al. (2024). As illustrated in the table, filtering with ICC outperforms by a large margin the alternative filtering methods for captioning given a fixed number of desired samples and training iterations. Note that unlike other methods, ICC is directly aligned with the captioning objective, as a captioning model should generate visually-grounded concrete text. This may explain the large gap in performance between ICC and other filtering baselines. We show qualitative comparison between captioning models trained with different filtering methods in Figure 4, exemplifying how filtering with ICC promotes more concrete and accurate captioning.